# Training Set Reconstruction from Differentially Private Forests: How Effective is DP?

### v1

@@ 5 Related Work
Reconstruction Attacks. Reconstruction attacks were first introduced against database queries by Dinur & Nissim (2003). In their setup, all database columns are public except for a binary one, which the adversary attempts to recover using (noisy) responses to counting queries. By modeling the noise within a linear programming framework, their attack successfully retrieves the private bit for all individuals after a significant number of queries. Over the following decade, several works extended this approach (Dwork et al., 2017).
@@
More recently, reconstruction attacks have been adapted to machine learning models. Unlike database-focused attacks, these methods often aim to infer specific attributes or recover information about individual samples rather than reconstructing the entire training set (Fredrikson et al., 2015). Many of these attacks also rely on auxiliary information beyond the trained model, such as intermediate gradients during training (Phong et al., 2017; Salem et al., 2020), stationary points reached during optimization (Haim et al., 2022), or fairness-related constraints (Hu & Lan, 2020; Hamman et al., 2022; Ferry et al., 2023).
@@
In contrast, recent studies have demonstrated that tree-based models are particularly vulnerable to reconstruction attacks due to their combinatorial structure. For example, Gambs et al. (2012) and Ferry et al. (2024a) developed algorithms leveraging white-box access to a single decision tree to reconstruct a probabilistic representation of its dataset, encoding all reconstructions consistent with the tree’s structure. More recently, Ferry et al. (2024b) proposed DRAFT, a CP-based reconstruction attack against RFs, showing that the structure of an RF can be exploited to reconstruct most of its training data efficiently.
@@
DP (Dwork & Roth, 2014) has been widely adopted as a protection mechanism as it provides strong theoretical guarantees. However, while it inherently bounds the risk of membership inference attacks, the effectiveness of its protection against other types of attacks varies. For instance, DP has been shown to mitigate reconstruction attacks on deep neural networks (Ziller et al., 2024) and language models (Stock et al., 2022), even for large privacy budgets. However, these defenses often come at a significant cost to model performance, and the considered attacks typically do not exploit the DP mechanism explicitly. Our work advances this field by explicitly accounting for the DP mechanism within the attack formulation. We analyze the trade-offs between the enforced DP budget, the empirical reconstruction attack success, and the DP model’s predictive performances in the context of tree ensembles.
@@
Building DP Random Forests. Traditional tree induction algorithms iteratively perform a greedy selection of the attributes and values to split on at each level of the considered tree, based on some information gain criterion. Some works on DP RFs proceed similarly (Patil & Singh, 2014; Fletcher & Islam, 2015b; Rana et al., 2015). However, since the training data need to be accessed at each level of the tree to greedily determine the splits, the DP budget $\varepsilon$ must be divided between them, resulting in the addition of a significant amount of noise. Thus, other approaches proposed to randomly determine all splitting attributes and values (Jagannathan et al., 2012; Fletcher & Islam, 2015a, 2017), and only access the data to populate the leaves. Even if random splits may partition the data imperfectly, this reduction in the amount of noise required to ensure DP generally allows for better predictive performance. Note that this specific aspect makes no difference in our attack’s design, since the reconstruction model does not directly exploit the way the splits are determined.
@@
In non-DP RFs, bootstrap sampling is commonly used to generate each tree’s training set, by performing sampling with replacement from $N$ to $N$. While some DP RF approaches (Rana et al., 2015) have adopted this strategy, providing tight and thorough DP guarantees remains challenging, as bootstrapped subsets are not disjoint and often contain multiple occurrences of certain examples. An alternative is to divide the original dataset into disjoint subsets to independently train each tree, enabling parallel composition of the privacy budget (Fletcher & Islam, 2015a, 2017). However, this approach may harm predictive performance as each tree is trained on significantly less data. Moreover, this also results in smaller leaves’ class counts, which in turn are proportionally more affected by the noise added to enforce DP. A standard approach is to learn all trees on the full training dataset, even though it requires dividing the privacy budget across trees (Jagannathan et al., 2012; Fletcher & Islam, 2015a, 2017; Patil & Singh, 2014; Fletcher & Islam, 2015b). While our attack could handle bootstrap sampling as in Ferry et al. (2024b), if the trees are trained on completely disjoint subsets, a per-tree probabilistic reconstruction approach, such as the one proposed by Ferry et al. (2024a), suffices, since nothing can be gained by relating the information discovered on examples across different trees.
@@
Finally, some approaches such as IBM’s diffprivlib implementation (Holohan et al., 2019), release only (noisy) majority labels within the leaves of DP RFs to conserve the privacy budget instead of providing full class counts. However, this strategy prevents the use of the traditional soft voting scheme for inference, relying instead on simple majority voting, which may degrade predictive performance. Furthermore, access to class counts is often necessary for auditing purposes. Adapting the attack developed in this study to handle this setup would require modifying the objective function and introducing additional variables, to compute the likelihood of generating an observed (noisy) majority label for each leaf based on the guessed class counts.
@@ 6 Conclusion
This work has demonstrated that even with a low privacy budget, accurate reconstruction of the training examples of an $\varepsilon$-DP RF remains possible. Crucially, it shows that a model can comply with rigorous DP guarantees while still leaking a significant portion of its training data, highlighting that privacy protection cannot be solely reduced to choosing a sufficiently small $\varepsilon$. Rather, the design of the DP mechanism itself plays a fundamental role in mitigating reconstruction risks. This aligns with recent findings on privacy attacks (i.e., re-identification and attribute inference) against local-differentially private protocols, in which the attack success depends not only on $\varepsilon$ but also on the mechanism’s structure (Arcolezi et al., 2023). Similarly, our results suggest that a careful selection of the hyperparameters used to tune the RF and the DP mechanism allows for better trade-offs between predictive performance and empirical protection against reconstruction attacks.
@@
Impact Statement
However, DP does not (at least directly) provide theoretical protection against dataset reconstruction attacks. Our work demonstrates that, in practice, machine learning models can comply with rigorous DP guarantees while still leaking a significant portion of their training data. This critical finding highlights the need for developing new defense mechanisms, and emphasizes the importance of the design of the DP mechanism itself, beyond the sole privacy budget. Going in this direction, our study provides insights into the impact of different DP design choices and advocates for empirically evaluating a model’s resilience against a variety of reconstruction attacks —including the one introduced in this work— before disclosure. Our findings emphasize that achieving privacy-preserving machine learning requires a holistic approach that carefully balances transparency, model utility, and empirical robustness against privacy threats. Ultimately, DP guarantees alone are not a silver bullet for privacy.
@@ A.1 Confidence Interval on $N$
While reconstruction attacks targeting a complete dataset recovery in the literature generally assume knowledge of its size $N$ (Dwork et al., 2017), our approach can be slightly modified to be applicable without this information. In this section, we provide and empirically evaluate a CP formulation derived from the one presented in Section 3 that does not rely on knowing $N$. Note that while this setup is both more generic and more realistic, it is also more challenging.
When $N$ is unknown, it becomes a variable to infer. To bound the search, we will first calculate a confidence interval $\llbracket N_{\textsc{min}},N_{\textsc{max}}\rrbracket$. According to the DP RF algorithm in Section 2.2, each training example is used exactly once per tree. For a tree $t\in\mathcal{T}$, the sum of the non-noisy leaves’ counts equals $N$. In our case, since the Laplace noise is centered at zero, for a tree $t\in\mathcal{T}$, we can expect the sum $N_{t}^{*}=\smash{\sum\limits_{v,c}n_{tvc}^{*}}$ to lie close to $N$.
@@
where $t_{95}$ is the Student’s t-coefficient for a 95% confidence level according to $\lvert\mathcal{T}\rvert$.
@@ A.2 Adapted Constraint Programming Formulation
As in the formulation presented in Section 3, for each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ of each tree $t\in\mathcal{T}$, we let variables $n_{tvc}\in\{\max(0,n_{tvc}^{*}-\delta),\dots,n_{tvc}^{*}+\delta\}$ and $\Delta_{tvc}\in\{-\delta,\dots,\delta\}$ respectively model the (guessed) number of examples of class $c$ within this leaf, and the amount of noise added to it. We also use variable $\smash{\widetilde{N}}\in\{N_{\textsc{min}},\dots,N_{\textsc{max}}\}$ encode the (guessed) number of reconstructed examples.
@@
Finally, we define a set of variables modeling the path of each example $k\in\{1,\dots,N_{\textsc{max}}\}$ through each tree $t\in\mathcal{T}$: variable $y_{tvkc}\in\{0,1\}$ indicates if it is classified in class $c\in\mathcal{C}$ by leaf $v\in\smash{\mathcal{V}_{t}^{L}}$. Given these variables, as previously, we define constraints to compute class counts and connect them to the assignments of examples to leaves:
@@
Constraint (10) ensures that the examples’ counts sum up to $\widetilde{N}$ within each tree. Constraint (11) guarantees that each example is associated to exactly one class. Constraint (12) enforces that each example can only contribute to the counts of its class. Note that this constraint differs from the formulation presented in Section 3. Indeed, if $z_{kc}=1$ but $k>\widetilde{N}$, example $k$ is actually not part of the reconstructed dataset, and hence is not assigned to any leaf of any tree, as stated in Constraint (13). Constraint (14) ensures consistency between the examples’ assignments to the leaves, and the (guessed) counts.
@@
and the original objective is retained with $\Delta_{tvc}=n_{tvc}^{*}-n_{tvc}$:
@@ A.3 Experimental Evaluation
To assess the effectiveness of our reconstruction attack in a context where the number of training examples $N$ is unknown, we reproduce the experimental setup as described in Section 4.1, using our modified attack formulation introduced in the previous subsection. The computation of the reconstruction error must be slightly adapted. When $N$ is known, the attack’s performance is assessed by determining the correspondence between reconstructed and original examples using a minimum-cost matching. The average distance between these matched pairs defines the reconstruction error. However, when $N$ is unknown, it is possible that the inferred reconstruction contains more (or less) than $N$ examples, making this pairwise matching not directly applicable. In the case where the number of reconstructed examples exceeds the actual one ($\widetilde{N}>N$), we randomly duplicate some (true) training examples. In contrast, if the number of reconstructed examples is smaller than the actual one ($\widetilde{N}<N$), we randomly subsample the same number of examples from the original training set. The same minimum-cost matching is then performed, and the average distance between pairs of matched examples defines the reconstruction error.
The results of these experiments are presented in Table 2 for all datasets, forest sizes $\lvert\mathcal{T}\rvert$, and tree depths $d$. Comparing these results with those in Table 1 reveals two key trends. First, a few runs failed to retrieve a feasible reconstruction within the given time frame ($2$ hours), indicating that the attack is computationally more expensive in this setup. This is expected, as the modified CP model introduces additional variables and constraints, leading to a larger search space and calling for longer solution times. Still, in most cases, the attack finds a feasible reconstruction. Second, and more importantly, the reconstruction error remains comparable to the case where $N$ is known. Despite the increased computational cost, the attack is still able to retrieve meaningful information encoded by the forest regarding its training data in most cases, thereby confirming our Result 5.
Table 2: Experimental results without assuming the training set size $N$ is known. The table reports the average reconstruction error across varying numbers of trees $\lvert\mathcal{T}\rvert$, tree depths $d$, and privacy budgets $\varepsilon$ for the three datasets. Cases where the solver did not find a feasible reconstruction within the $2$-hour time limit are indicated by “–”.
@@ Appendix B Determining a Search Interval for Noise Values
Rather than modeling all possible (integer) random noise values $Y_{tvc}$ within $\mathbb{Z}$, our CP model focuses on a (sufficient) predefined range $\{-\delta,\dots,\delta\}$ such that $\mathbb{P}(\textrm{int}(Y_{tvc})\in\{-\delta,\dots,\delta\})\geq 0.999$. In practice, we use $\delta=\lceil 12/\varepsilon_{v}\rceil$. As displayed in Figure 5, this value is suitable for all considered privacy budgets $\varepsilon_{v}$. Indeed, it always leads to the definition of slightly wider search intervals for the variables modelling the noise, theoretically guaranteeing that the probability of the actual noise value being within it is greater than $0.999$.
@@
Appendix C Trade-offs between Predictive Performance and Reconstruction Error – Detailed Results

### v2

@@ 5 Related Work
Reconstruction Attacks and DP. Reconstruction attacks were first introduced against database queries by Dinur and Nissim [2003]. In their setup, all database columns are public except for a binary one, which the adversary attempts to recover using (noisy) responses to counting queries. By modeling the noise within a linear programming framework, their attack successfully retrieves the private bit for all individuals. Over the following decade, several works extended this approach [Dwork et al., 2017]. More recently, reconstruction attacks have been adapted to machine learning models. Unlike database-focused ones, these attacks often aim to infer specific attributes or recover information about individual samples rather than reconstructing the entire training set [Fredrikson et al., 2015]. Many of these attacks also rely on auxiliary information beyond the trained model, such as intermediate gradients computed during training [Phong et al., 2017, Salem et al., 2020], stationary points reached during optimization [Haim et al., 2022] or fairness-related constraints [Hu and Lan, 2020, Hamman et al., 2022, Ferry et al., 2023].
@@
In contrast, recent studies have shown that tree-based models are vulnerable to reconstruction attacks due to their combinatorial structure. For example, Gambs et al. [2012] and Ferry et al. [2024a] developed algorithms leveraging white-box access to a single decision tree to reconstruct a probabilistic representation of its dataset, encoding all reconstructions consistent with the tree’s structure. More recently, Ferry et al. [2024b] proposed DRAFT, a CP-based reconstruction attack against RFs, showing that the structure of an RF can be exploited to reconstruct most of its training data efficiently.
@@
DP has been shown empirically to mitigate reconstruction attacks on deep neural networks [Ziller et al., 2024] and language models [Stock et al., 2022]. However, these defenses often come at a significant cost to model performance, and the considered attacks typically do not exploit the DP mechanism explicitly. Our work advances this field by (1) explicitly accounting for the DP mechanism within the attack formulation, and (2) analyzing the trade-offs between the DP budget, the empirical reconstruction attack success and the DP model’s predictive performances in the context of tree ensembles.
@@
Building DP Random Forests. Traditional tree induction algorithms iteratively perform a greedy selection of the attributes and values to split on at each level of the considered tree, based on some information gain criterion. Some works on DP RFs proceed similarly [Patil and Singh, 2014, Fletcher and Islam, 2015b, Rana et al., 2015]. However, since the training data needs to be accessed at each level of the tree to greedily determine the splits, the DP budget $\varepsilon$ must be divided between them, resulting in the addition of a significant amount of noise. Thus, other approaches proposed to randomly determine all splitting attributes and values [Jagannathan et al., 2012, Fletcher and Islam, 2015a, 2017], and only access the data to populate the leaves. Even if random splits may partition the data imperfectly, this reduction in the amount of noise required to ensure DP generally allows for better predictive performance. Note that this specific aspect makes no difference in our attack’s design, since the reconstruction model does not directly exploit the way the splits are determined.
@@
In non-DP RFs, bootstrap sampling (bagging) is commonly used to generate each tree’s training set, by performing sampling with replacement from $N$ to $N$. While some DP RF approaches [Rana et al., 2015] have adopted this strategy, providing tight and thorough DP guarantees remains challenging, as bootstrapped subsets are not disjoint and often contain multiple occurrences of certain examples. An alternative is to divide the original dataset into disjoint subsets to independently train each tree, enabling parallel composition of the privacy budget [Fletcher and Islam, 2015a, 2017]. However, this approach may harm predictive performance as each tree is trained on significantly less data. Moreover, this also results in smaller leaves’ class counts, which in turn are proportionally more affected by the noise added to enforce DP. A standard approach is to learn all trees on the full training dataset, even though it requires dividing the privacy budget across trees [Jagannathan et al., 2012, Fletcher and Islam, 2015a, 2017, Patil and Singh, 2014, Fletcher and Islam, 2015b]. Our attack could handle bagging as in Ferry et al. [2024b]. However, if the trees are trained on disjoint subsets, a per-tree probabilistic reconstruction approach such as the one proposed by Ferry et al. [2024a] suffices, since nothing can be gained by relating the information discovered across different trees.
@@
Finally, some approaches such as IBM’s diffprivlib implementation [Holohan et al., 2019], release only (noisy) majority labels within the leaves of DP RFs to conserve the privacy budget instead of providing full class counts. However, this strategy prevents the use of the traditional soft voting scheme for inference, relying instead on simple majority voting, which may degrade predictive performance. Furthermore, the access to class counts is often necessary for auditing purposes. Adapting the attack developed in this study to handle this setup would require modifying the objective function and introducing additional variables, to compute the likelihood of generating an observed (noisy) majority label for each leaf based on the guessed class counts.
@@ 6 Conclusion
This work has demonstrated that even with a low privacy budget, accurate reconstruction of the training examples of an $\varepsilon$-DP RF remains possible. Crucially, it shows that a model can comply with rigorous DP guarantees while still providing significant information regarding its training dataset to potential reconstruction attackers. This highlights that privacy protection cannot be solely reduced to choosing a sufficiently small $\varepsilon$. Rather, the design of the DP mechanism itself plays a fundamental role in mitigating reconstruction risks. This aligns with recent findings on privacy attacks (i.e., re-identification and attribute inference) against local-differentially private protocols, in which the attack success depends not only on $\varepsilon$ but also on the mechanism’s structure [Arcolezi et al., 2023]. Similarly, our results suggest that a careful selection of the hyperparameters used to tune the RF and the DP mechanism allows for better trade-offs between predictive performance and empirical protection against reconstruction attacks.
@@
Acknowledgements
This research was enabled by support provided by Calcul Québec and the Digital Research Alliance of Canada, as well as funding from the SCALE AI Chair in Data-Driven Supply Chains.
@@ A.1 Confidence Interval on $N$
While reconstruction attacks targeting a complete dataset recovery in the literature generally assume knowledge of its size $N$ [Dwork et al., 2017], our approach can be slightly adapted to be applicable without this information. In this section, we provide and empirically evaluate a CP formulation derived from the one presented in Section 3 that does not rely on knowing $N$. While this setup is both more generic and more realistic, it is also more challenging.
When $N$ is unknown, it becomes a variable to infer. To bound the search, we will first calculate a confidence interval $\llbracket N_{\textsc{min}},N_{\textsc{max}}\rrbracket$. According to the DP RF algorithm in Section 2.2, each training example is used exactly once per tree. Thus, for a tree $t\in\mathcal{T}$, the sum of the non-noisy leaves’ counts equals $N$. In our case, since the Laplace noise is centered at zero, for a tree $t\in\mathcal{T}$, we can expect the sum $N_{t}^{*}=\smash{\sum\limits_{v,c}n_{tvc}^{*}}$ to lie close to $N$.
@@
in which $t_{95}$ is the Student’s t-coefficient for a 95% confidence level according to $\lvert\mathcal{T}\rvert$.
@@ A.2 Adapted Constraint Programming Formulation
As in the formulation presented in Section 3, for each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ of each tree $t\in\mathcal{T}$, we let variables $n_{tvc}\in\{\max(0,n_{tvc}^{*}-\delta),\dots,n_{tvc}^{*}+\delta\}$ and $\Delta_{tvc}\in\{-\delta,\dots,\delta\}$ respectively model the (guessed) number of examples of class $c$ within this leaf and the amount of noise added to it. We also use variable $\smash{\widetilde{N}}\in\{N_{\textsc{min}},\dots,N_{\textsc{max}}\}$ to encode the (guessed) number of reconstructed examples.
@@
Finally, we define a set of variables modelling the path of each example $k\in\{1,\dots,N_{\textsc{max}}\}$ through each tree $t\in\mathcal{T}$: variable $y_{tvkc}\in\{0,1\}$ indicates if it is classified in class $c\in\mathcal{C}$ by leaf $v\in\smash{\mathcal{V}_{t}^{L}}$. Given these variables, as previously, we define constraints to compute class counts and connect them to the assignments of examples to leaves:
@@
Constraint (10) ensures that the examples’ counts sum up to $\widetilde{N}$ within each tree. Constraint (11) guarantees that each example is associated with exactly one class. Constraint (12) enforces that each example can only contribute to the counts of its class. Note that this constraint differs from the formulation presented in Section 3. Indeed, if $z_{kc}=1$ but $k>\widetilde{N}$, example $k$ is actually not part of the reconstructed dataset and hence is not assigned to any leaf of any tree, as stated in Constraint (13). Constraint (14) ensures consistency between the examples’ assignments to the leaves and the (guessed) counts.
@@
The original objective is also retained with $\Delta_{tvc}=n_{tvc}^{*}-n_{tvc}$:
@@ A.3 Experimental Evaluation
To assess the effectiveness of our reconstruction attack in a context where the number of training examples $N$ is unknown, we reproduce the experimental setup as described in Section 4.1, using our modified attack formulation introduced in the previous subsection. The computation of the reconstruction error must be slightly adapted. Indeed when $N$ is known, the attack’s performance is assessed by determining the correspondence between reconstructed and original examples using a minimum-cost matching. The average distance between these matched pairs defines the reconstruction error. However, when $N$ is unknown, it is possible that the inferred reconstruction contains more (or less) than $N$ examples, making this pairwise matching not directly applicable. In the case in which the number of reconstructed examples exceeds the actual one ($\widetilde{N}>N$), we randomly duplicate some (true) training examples. In contrast, if the number of reconstructed examples is smaller than the actual one ($\widetilde{N}<N$), we randomly subsample the same number of examples from the original training set. The same minimum-cost matching is then performed and the average distance between pairs of matched examples defines the reconstruction error.
The results of these experiments are presented in Table 2 for all datasets, forest sizes $\lvert\mathcal{T}\rvert$ and tree depths $d$. Comparing these results with those in Table 1 reveals two key trends. First, a few more runs failed to retrieve a feasible reconstruction within the given time frame ($2$ hours), indicating that the attack is computationally more expensive in this setup. This is expected, as the modified CP model introduces additional variables and constraints, leading to a larger search space and calling for longer solution times. Still, in most cases, the attack finds a feasible reconstruction. Second, and more importantly, the reconstruction error remains comparable to the case in which $N$ is known. Despite the increased computational cost, the attack is still able to retrieve meaningful information encoded by the forest regarding its training data in most cases, thereby confirming our Result 5.
Table 2: Experimental results without assuming the training set size $N$ is known. The table reports the average reconstruction error across varying numbers of trees $\lvert\mathcal{T}\rvert$, tree depths $d$ and privacy budgets $\varepsilon$ for the three datasets. The cases for which the solver did not find a feasible reconstruction within the $2$-hour time limit are indicated by “–”.
@@ Appendix B Determining a Search Interval for Noise Values
Rather than modelling all possible (integer) random noise values $Y_{tvc}$ within $\mathbb{Z}$, our CP model focuses on a (sufficient) predefined range $\{-\delta,\dots,\delta\}$ such that $\mathbb{P}(\textrm{int}(Y_{tvc})\in\{-\delta,\dots,\delta\})\geq 0.999$. In practice, we use $\delta=\lceil 12/\varepsilon_{v}\rceil$. As displayed in Figure 5, this value is suitable for all considered privacy budgets $\varepsilon_{v}$. Indeed, it always leads to the definition of slightly wider search intervals for the variables modelling the noise, theoretically guaranteeing that the probability of the actual noise value being within it is greater than $0.999$.
@@
Appendix C Additional Experimental Results
C.1 Running Times
The average running time (over all performed experiments) to find a feasible reconstruction is below two minutes. Table 3 details the running time required by the CP solver to find a feasible reconstruction, averaged over the 5 random seeds for all considered datasets, numbers of trees $\lvert\mathcal{T}\rvert$ within the target forest, tree depths $d$ and privacy budgets $\varepsilon$. Several general trends can be observed. Increasing the number of trees or their depth leads to larger CP models, including more variables, more constraints and larger search spaces. This logically increases the solver’s running time. Considering tight privacy budgets also widens the solver’s search space and running time. Nevertheless, in all cases, the CP solver is able to find a feasible reconstruction using only a fraction of its predefined 2-hours timeout, and can then use the remaining time to refine it and improve its objective function (corresponding to the noise likelihood, as explained in Section 3).
Table 3: Average running time (s) to get a feasible reconstruction for different numbers of trees $\lvert\mathcal{T}\rvert$ within the target forest, tree depths $d$ and privacy budgets $\varepsilon$ for the three datasets. The situation in which the solver failed to find a feasible reconstruction within the $2$-hour time limit are marked with “–”.
Adult COMPAS Default Credit
$d=3$ $d=5$ $d=7$ $d=3$ $d=5$ $d=7$ $d=3$ $d=5$ $d=7$
$\lvert\mathcal{T}\rvert=1$ $\varepsilon=0.1$ 1.81 9.73 24.98 1.95 7.10 22.77 2.65 9.48 28.37
$\varepsilon=1$ 0.56 3.90 16.57 0.51 2.75 10.68 0.56 3.63 15.53
$\varepsilon=5$ 0.45 3.15 13.57 0.49 2.61 9.89 0.53 3.60 14.84
$\varepsilon=10$ 0.47 3.16 13.78 0.47 2.65 10.03 0.51 3.69 16.69
$\varepsilon=20$ 0.46 1.95 13.35 0.39 2.26 9.66 0.37 1.65 12.21
$\varepsilon=30$ 0.41 1.57 12.69 0.43 2.50 9.74 0.48 2.78 15.54
$\lvert\mathcal{T}\rvert=5$ $\varepsilon=0.1$ 14.35 47.58 123.09 10.41 39.33 96.10 12.94 41.43 128.34
$\varepsilon=1$ 6.61 22.23 60.73 4.84 24.05 67.07 4.52 23.36 56.40
$\varepsilon=5$ 3.72 23.87 53.88 3.72 25.14 71.13 3.56 28.15 79.50
$\varepsilon=10$ 4.23 26.25 75.20 3.82 22.80 73.11 3.00 24.55 88.45
$\varepsilon=20$ 2.95 24.62 85.41 3.32 20.51 63.55 3.82 22.55 54.54
$\varepsilon=30$ 2.68 24.47 110.77 3.60 27.20 51.81 3.27 22.18 118.82
$\lvert\mathcal{T}\rvert=10$ $\varepsilon=0.1$ 34.90 94.27 284.07 27.20 75.29 422.43 31.08 108.73 442.63
$\varepsilon=1$ 17.79 42.57 157.09 18.73 45.19 103.73 14.70 48.83 109.73
$\varepsilon=5$ 10.88 43.60 93.64 9.19 41.73 83.29 12.10 44.93 97.78
$\varepsilon=10$ 10.88 47.81 98.43 10.72 44.00 125.38 12.11 100.56 122.64
$\varepsilon=20$ 10.35 40.89 106.39 8.62 39.82 119.52 12.33 63.34 124.82
$\varepsilon=30$ 8.68 46.77 90.21 12.32 46.17 87.92 9.17 41.91 284.96
$\lvert\mathcal{T}\rvert=20$ $\varepsilon=0.1$ 90.69 293.24 1033.74 84.66 282.49 968.48 97.05 261.93 1022.18
$\varepsilon=1$ 32.68 92.48 357.22 26.43 89.08 300.26 30.97 85.97 354.49
$\varepsilon=5$ 22.18 65.19 167.35 22.18 68.28 206.02 17.92 66.48 225.78
$\varepsilon=10$ 22.39 68.52 234.35 27.08 67.73 254.61 19.43 71.50 242.74
$\varepsilon=20$ 28.65 225.42 360.96 25.57 125.94 279.20 21.58 243.11 890.48
$\varepsilon=30$ 35.18 233.34 387.37 26.55 72.33 285.36 25.58 127.86 511.48
$\lvert\mathcal{T}\rvert=30$ $\varepsilon=0.1$ 185.31 842.61 - 170.20 699.23 - 150.73 752.15 -
$\varepsilon=1$ 47.45 152.51 611.71 39.69 146.58 666.69 41.66 158.16 451.77
$\varepsilon=5$ 29.43 101.53 405.87 26.72 95.50 259.35 31.40 89.66 438.50
$\varepsilon=10$ 31.83 92.64 363.51 28.17 86.97 333.50 30.04 96.03 380.74
$\varepsilon=20$ 34.24 87.08 357.17 58.27 212.95 375.59 32.66 119.71 413.38
$\varepsilon=30$ 45.91 410.98 536.70 32.32 217.46 828.98 36.58 232.30 862.32
C.2 Trade-offs between Predictive Performance and Reconstruction Error